# SemHiTok: A Unified Image Tokenizer via Semantic-Guided Hierarchical Codebook for Multimodal Understanding and Generation

### v4

@@ 4.1 Experimental Setup
Table 3: Training details of tokenizer.
Settings Semantic Codebook Training Pixel Reconstruction Enablement
Stage1 Stage2 Enhance Decoder
Learning rate 1e-4 1e-4 4e-5 2e-5
Batch size 256 256 256 256
Data 50M COYO ImageNet-1K 20M COYO 20M COYO&20M MJ
Training Module Semantic Modules Pixel Modules Pixel Modules Pixel Decoder
Tokenizer. For semantic branch, we employ SigLIP[66](SigLIP-large-p16 and SigLIP-so400m-patch14-384) as the semantic encoder and three self-attention layers as semantic decoder to reconstruct semantic features from SigLIP. In addition, our semantic codebook size is 16384, which updated using EMA[17] algorithm to enhance stability. For pixel branch, We employ ViT as both the pixel encoder and decoder, assigning 8 pixel sub-codes to each semantic code. The pixel sub-codebooks use normalized VQ structure which enhances both usage and stability. The training of SemHiTok is conducted in two stages. During semantic codebook training, we train the semantic tokenizer for one epoch on 50M subset of COYO-700M[4]. For PRE stage, we first train the model(ViT-Base) on ImageNet[10], and then fine-tune on 50M COYO to improve its generalization, following LlamaGen[42]. To further improve the reconstruction and generation performance, we enlarge the size of the pixel decoder(ViT-Large) and fine-tune on the 20M COYO data and 20M MidJourney-style synthetic data. More tokenizer detail is shown in Tab.3.
Unified MLLM. We use Qwen2.5[60] 7B as base LLM, and expand it’s vocabulary and output head layer. We evaluate visual understanding on standard VQA benchmarks including SEEDB[21], POPE[26], GQA[16], MMMU[64], MMB[32] and MME[12]. For visual generation evaluation, we report results on MJHQ-30K[23] and GenAI-Bench[20]. Follow existing work[54; 34], we first pretrain the model and adapter layer on a mix of multimodal data, which is composed of 3.5M language data from Magpie[59] and Openorca[27], 10M caption image-text pairs data and 12M MidJourney-style synthetic data. Subsequently, we finetune the model on 1M language dataset from Magpie[59] and Evol-Instruct[58], 4M MidJourney-style synthetic data and 4M understanding data from emova[7] and LLaVA-SFT[31].
@@ 4.2 Unified Image Tokenizer
Table 4: Comparative analysis of tokenizers on multimodal comprehension tasks. *: Both tokenizer and LLM are reproduced in our setting. Our method achieves SOTA performance compare with other discrete tokenizer.
@@
SigLIP[66] Vicuna-7B LLaVA-v1.5 256 83.76 1481.0 65.28 61.9
LlamaGen[61] Vicuna-7B LLaVA-v1.5 256 65.6 716.8 35.0 39.8
VILA-U[54] Vicuna-7B LLaVA-v1.5 256 81.6 1311.6 56.9 55.3
UniTok[34] Vicuna-7B LLaVA-v1.5 256 81.7 1303.7 56.6 61.1
TokLIP[28] Qwen2.5-7B-Ins LLaVA-v1.5 256 81.2 1346.8 59.8 57.4
SDE*[57] Vicuna-7B LLaVA-v1.5 256 77.3 1240.0 56.7 58.0
@@
Image Reconstruction. We present reconstruction performance of SemHiTok on ImageNet-50k validation set in Tab.2. Notably, SemHiTok excels in reconstruction quality compared to unified tokenizer, recording an impressive 1.16 rFID with 16×16$\times$16 × downsampling ratio. Furthermore, our tokenizer is competitive with some domain-specific tokenizer such as IBQ and FQGAN, while significantly outperforming previous methods RQVAE and VQGAN-LC. At a resolution of 384, our tokenizer further improves the rFID score to 0.66. The results validate the effectiveness of SGHC design in modeling pixel feature space of semantic code.
LLaVa-v1.5 Multimodal Understanding. To ensure a fair comparison, we conduct experiments to evaluate the multimodal understanding performance of existing open-source tokenizers and SemHiTok under the standard LLaVA-v1.5[31] setting. The results are as shown in Tab.4. Due to the lack of pre-alignment with text, LlamaGen performs the worst performance. Due to hybrid structure or joint training strategies, there remains a significant gap between previous tokenizer and continuous representations in terms of understanding. However, SemHiTok achieves the SOTA of the discrete tokenizer, which is closest to the SigLIP of the continuous input.
The above two groups of experiments show that SemHiTok can extract unified image features, which includes low-level pixel features and high-level semantic features, and has the potential to be applied to unified MLLM.
4.3 Unified MLLM
Multimodal Understanding. We evaluate the understanding performance of SemHiTok on diverse benchmarks in Table 5. Compared to VILA-U(256), our model achieves performance gains of 13.4 and 112.8 points respectively on SEED and MME-P benchmark. Notably, our model surpasses expert-level models on key benchmarks, achieving 2.1 and 3.7 points higher than ShareGPT4V on MMMU and MMB, respectively. This bridges the gap between discrete visual tokens and continuous visual tokens in multimodal understanding tasks, strongly demonstrating the superiority and potential of our approach. With a higher-resolution version of SemHiTok, the model’s performance is further improved. Among Und&Gen. Discrete, SemHiTok achieves state-of-the-art performance on most metrics, such as SEED, MMMU,MMB,MME and MME-P. Visualizations on understanding tasks are shown in Fig.7.
Figure 7: Visualizations on understanding tasks.
Table 5: Quantitative results on multimodal understanding benchmarks. SemHiTok achieves SOTA performance on most benchmarks among Und&Gen Discrete MLLMs, and is comparable to or even surpasses some Only Und and Und&Gen. Continuous models. The performance on Und&Gen Discrete with top-1 an top-2 value are denoted in bold and underline respectively.
@@ 4.3 Unified MLLM
LLaVA-Phi[69] 2.7B 256 - 85.0 - - 59.8 - 1335.1 28.9
LLaVA-v1.5[31] 7B 336 58.6 85.9 62.0 35.4 64.3 - 1510.7 31.1
Qwen-Vl-Chat[60] 7B 448 57.7 - 57.5 30.5 - 1848.3 1487.5 -
ShareGPT4V[8] 7B 336 69.7 - 63.3 37.2 68.8 1943.8 1567.4 37.6
@@
Unified-IO 2[33] 6.8B 384 - 87.7 - - - - - -
DreamLLM[11] 7B 224 - - - - - - - 26.6
LaVIT[18] 7B 224 - - 46.8 - 58.0 - - -
Janus[52] 1.3B 384 63.7 87.0 59.1 30.5 69.4 - 1338.0 34.3
@@
LWM[29] 7B 256 - 75.2 44.8 - - - - 9.6
SEED-LLaMA[22] 13B 256 53.7 - - - - - - -
Show-o[56] 1.5B 256 - 80.0 - 26.7 - - 1097.2 -
Liquid[53] 7B 512 - 81.1 71.3 - - - 1119.3 -
EMU3[50] 8B 512 68.2 85.2 60.3 31.6 58.5 1509.9 1243.8 37.2
VILA-U[54] 7B 256 56.3 83.9 58.3 - - - 1336.2 27.7
VILA-U[54] 7B 384 59.0 85.8 60.8 - - - 1401.8 33.5
UniTok[34] 7B 256 - 84.0 61.1 - - - 1448.0 -
@@
Figure 8: Presentation of the generated results.
Text-to-Image Generation. To evaluate the text-to-image generation, we use GenAI-bench[20] and MJHQ30K[23] benchmark, and the results are shown in Tab8. For GenAI-bench, we use clip-flant5-xxl as VQA score model to reflect the consistency between text descriptions and generated images. On this challenging benchmark, our model achieves competitive performance, with advanced score only lower than Liquid, while outperforming all other unified models. Furthermore, our model even outperformed some diffusion-based expert models, such as SDXL and SD v2. The strong results underscore the superior capability of our unified MLLM in complex text-to-image generation tasks. For MJHQ30K[23], we use the generation FID metirc on generatede images and high-qulity images. On this benchmark, SemHiTok-256 attains 5.40 gFID, setting a new state-of-the-art in autoregressive image generation. In addition, the generated images by our model is illustrated in Fig. 8, using prompts from the MJHQ30K dataset.
Table 6: Comparison of generation quality on GenAI [20] and MJHQ30K [23]. SemHiTok achieves comparable results with specialist models and unified MLLMs. The performance on Und&Gen. with top-1 an top-2 value are denoted in bold and underline respectively.
1 Model Params Type Res. GenAI-Bench MJHQ30K
Basic $\uparrow$ Advanced $\uparrow$ gFID $\downarrow$
@@
SD v2.1[39] Diffusion 1024 0.78 0.62
DALL-E 3[3] Diffusion 1024 0.90 0.70
PixArt-$\alpha$ [6] 0.6B Diffusion 1024 6.14
SDXL [35] 2.6B Diffusion 1024 0.83 0.63 9.55
Playgroundv2.5 [23] Diffusion 1024 4.48
@@
LWM [29] 7B Autoregressive 256 0.63 0.53 17.77
Show-o [56] 1.5B Discrete Diffusion 256 0.70 0.60 15.18
Janus [52] 1.3B Autoregressive 384 10.10
VILA-U [54] 7B Autoregressive 256 0.76 0.64 12.81
VILA-U [54] 7B Autoregressive 384 0.73 0.61 7.69
ILLUME [54] 7B Autoregressive 512 0.75 0.60 7.76
Liquid [53] 7B Autoregressive 512 0.83 0.65 5.47
UniTok [34] 7B Autoregressive 256 0.85 0.67 7.46
SemHiTok(ours) 7B Autoregressive 256 0.83 0.64 5.40
SemHiTok(ours) 7B Autoregressive 384 0.83 0.66 5.70
4.4 Ablation
Impact of Key Design. In Tab.7, we validate the impact of our key design choices in SemHiTok: semanitc codebook, semantic-guided hierarchical codebook(SGHC), Dual MLP and Enhance Deocoder. For efficiency, we only tested the understanding performance under LLaVA-v1.5, and train 40 epochs on ImageNet-1K for reconstruction tasks. We begin with the semantic codebook, which suffices for multimodal understanding but suffers from poor reconstruction. Incorporating SGHC enables pixel reconstruction and reduces rFID by 1.75, without noticeably affecting understanding performance. Introducing Dual-MLP further enhances multimodal understanding, even surpassing the semantic codebook alone, highlighting the effectiveness of multi-level feature modeling. Finally, a stronger pixel decoder brings additional improvements in reconstruction quality. Fig. 6 demonstrates the effects of various modules on reconstruction results. Compared to the semantic codebook, SGHC delivers more detailed reconstructions. Additionally, the enhanced decoder boosts performance on difficult samples.
Table 7: Impact of key design choices on reconstruction quality and multimodal understanding benchmarks. The gray bar represents the default setting in our experiments.
@@ 4.4 Ablation
Table 8: Impact of Concat type and sub-codebook size. w/o sem: not use semantic discrete token. The gray bar represents the default setting in our experiments.
@@
w/o 12 1.99 95.4%
Len 12 1.45 94.1%
Dim 8 1.42 96.4%
Dim 12 1.26 93.7%
Dim 16 1.19 79.3%
Table 9: Impact of VQ Type and dim of semantic codebook. We evaluate multimodal understanding performance under LLaVA-v1.5 setting. The gray bar represents the default setting in our experiments.
[…]
Norm 48 1249.6 52.2 52.8 34.8
Vanilla 48 1319.9 56.3 57.2 33.1
EMA 32 1385.9 58.3 61.2 35.1
EMA 48 1387.5 61.3 62.3 35.6
EMA 64 1428.7 60.9 62.5 35.5
Impact of Concat Type and Sub-Codebook Size. For efficiency, we only conduct training and evaluation on ImageNet-1K. We investigate the impact of concat type of semantic between pixel and the sub-codebook size, on reconstruction performance as shown in Tab. 9. For Concat type, w/o semantic token leads to significant drop in reconstruction quality, with s core increase of 0.73 compared to default setting. Furthermore, concatenation along the sequence length performs worse than along the dimension, as the two token sets are spatially aligned, making dimensional concatenation more appropriate. For sub-codebook size, increasing the size can improve the model’s reconstruction performance, but it exhibits marginal utility. In addition, the codebook usage significantly decreases when the sub-codebook size is set to 16, which indicates that too large a sub-codebook size is not cost-effective.
Impact of Semanitc VQ Type and Codebook dim. In Tab.9, we ablate VQ type and codebook dimension on multimodal understanding. Specifically, Norm VQ is not suitable for semantic discretization and shows the worst performance in understanding tasks. This indicates that it is difficult to discretely model complex and rich semantic information by Norm VQ. Replacing it with vanilla VQ brings a clear improvement. To further stabilize the semantic codebook, we use EMA VQ, which achieves the best results. For the semantic codebook dimension, higher values offer better representation but with marginal effect. We empirically set the dimension to 48 as the default.

### v5

@@ 3.1 Experimental Setup
Tokenizer. For the semantic branch, we employ SigLIP (Zhai et al., 2023) as the semantic encoder and three self-attention layers as the semantic decoder to reconstruct semantic features. For the pixel branch, we employ ViT as both the pixel encoder and decoder, assigning 8 pixel sub-codes to each semantic code. More tokenizer detail, please refer to Supplementary Material 7.3.
Unified MLLM. We use Qwen2.5-7B-Instruct (Yang et al., 2024) as the base LLM, and expand its vocabulary and output head layer. We evaluate visual understanding on standard VQA benchmarks including SEEDB (Li et al., 2023a), POPE (Li et al., 2023c), GQA (Hudson and Manning, 2019), MMMU (Yue et al., 2024), MMB (Liu et al., 2024b) and MME (Fu et al., 2023). For visual generation evaluation, we report results on MJHQ-30K (Li et al., 2024c), GenAI-Bench (Li et al., 2024a), GenEval (Ghosh et al., 2023) and DPG (Hu et al., 2024). More unified MLLM detail, please refer to Supplementary Material 7.4.
@@ 3.2 Unified Image Tokenizer
Table 2: Comparative analysis of tokenizers on multimodal comprehension tasks. *: Both the tokenizer and LLM are reproduced in our setting. Our method achieves SOTA performance compared with other discrete tokenizers.
@@
SigLIP (Zhai et al., 2023) Vicuna-7B LLaVA-v1.5 256 83.76 1481.0 65.28 61.9
LlamaGen (Sun et al., 2024a) Vicuna-7B LLaVA-v1.5 256 65.6 716.8 35.0 39.8
VILA-U (Wu et al., 2024c) Vicuna-7B LLaVA-v1.5 256 81.6 1311.6 56.9 55.3
SDE(MUSE-VL)* (Xie et al., 2025) Vicuna-7B LLaVA-v1.5 256 77.3 1240.0 56.7 58.0
TokLIP (Lin et al., 2025) Qwen2.5-7B-Ins LLaVA-v1.5 256 81.2 1346.8 59.8 57.4
@@
TokenFlow-384 (Qu et al., 2024) Vicuna-7B LLaVA-v1.5 384 84.9 1416.4 62.7 61.2
TokLIP (Lin et al., 2025) Qwen2.5-7B-Ins LLaVA-v1.5 384 82.7 1410.2 65.2 59.3
SemHiTok-384(Ours) Vicuna-7B LLaVA-v1.5 384 86.3 1465.6 64.1 62.3
Table 3: Quantitative results on multimodal understanding benchmarks. SemHiTok achieves SOTA performance on most benchmarks among Und&Gen Discrete MLLMs, and is comparable to or even surpasses some Only Und and Und&Gen. Continuous models. The performance on Und&Gen Discrete with top-1 and top-2 values is denoted in bold and underline, respectively.
@@
LLaVA-Phi (Zhu et al., 2024b) 2.7B 256 - 85.0 - - 59.8 - 1335.1 28.9
LLaVA-v1.5 (Liu et al., 2023b) 7B 336 58.6 85.9 62.0 35.4 64.3 - 1510.7 31.1
Qwen-Vl-Chat (Bai et al., 2023b) 7B 448 57.7 - 57.5 30.5 - 1848.3 1487.5 -
ShareGPT4V (Chen et al., 2024b) 7B 336 69.7 - 63.3 37.2 68.8 1943.8 1567.4 37.6
@@
LaVIT (Jin et al., 2023) 7B 224 - - 46.8 - 58.0 - - -
Janus (Wu et al., 2024a) 1.5B 384 63.7 87.0 59.1 30.5 69.4 - 1338.0 34.3
Janus-Pro-1B (Chen et al., 2025) 1.5B 384 68.3 86.2 58.9 38.9 65.5 - 1444.0 -
MAR (Wu et al., 2025) 1.5B 512 67.1 87.6 58.9 38.9 65.5 - 1155.0 -
@@
LWM (Liu et al., 2024a) 7B 256 - 75.2 44.8 - - - - 9.6
SEED-LLaMA (Li et al., 2024b) 13B 256 53.7 - - - - - - -
Show-o (Xie et al., 2024) 1.5B 256 - 80.0 - 26.7 - - 1097.2 -
Liquid (Wu et al., 2024b) 7B 512 - 81.1 71.3 - - - 1119.3 -
EMU3 (Wang et al., 2024c) 8B 512 68.2 85.2 60.3 31.6 58.5 1509.9 1243.8 37.2
VILA-U (Wu et al., 2024c) 7B 256 56.3 83.9 58.3 - - - 1336.2 27.7
VILA-U (Wu et al., 2024c) 7B 384 59.0 85.8 60.8 - - - 1401.8 33.5
UniToken (Jiao et al., 2025) 7B 384 69.3 - - 32.8 69.9 - - -
TokLIP (Lin et al., 2025) 7B 384 76.9 84.1 59.5 43.1 67.6 - 1488.4 29.8
TokenFlow-B (Qu et al., 2024) 13B 224 60.4 84.0 59.3 34.2 55.3 1660.4 1353.6 22.4
TokenFlow-L (Qu et al., 2024) 13B 256 62.6 85.0 60.3 34.4 60.3 1622.9 1365.4 27.7
SynerGen-VL (Li et al., 2025) 2.4B 512 62.0 85.3 59.7 34.2 53.7 1837.0 1381.0 34.5
@@
Table 4: Comparison of generation quality on GenAI and MJHQ30K. SemHiTok achieves comparable results with specialist models and unified MLLMs.
Model Params Type #Training Res. GenAI-Bench MJHQ30K
Images Basic $\uparrow$ Advanced $\uparrow$ gFID $\downarrow$
@@
SD v2.1 (Rombach et al., 2022) Diff 2000M 1024 0.78 0.62
DALL-E 3 (Betker et al., 2023) Diff - 1024 0.90 0.70
PixArt-$\alpha$ (Chen et al., 2023) 0.6B Diff - 1024 6.14
SDXL (Podell et al., 2023) 2.6B Diff 2000M 1024 0.83 0.63 9.55
Playgroundv2.5 (Li et al., 2024c) Diff - 1024 4.48
@@
LWM (Liu et al., 2024a) 7B AR - 256 0.63 0.53 17.77
Show-o (Xie et al., 2024) 1.5B Diff 36M 256 0.70 0.60 15.18
Janus (Wu et al., 2024a) 1.3B AR 65M 384 10.10
VILA-U (Wu et al., 2024c) 7B AR 15M 256 0.76 0.64 12.81
VILA-U (Wu et al., 2024c) 7B AR 15M 384 0.73 0.61 7.69
SynerGen-VL (Li et al., 2025) 2.4B AR 25M 256 - - 6.10
SDE(MUSE-VL) (Xie et al., 2025) 7B AR 10M 256 - - 7.73
ILLUME (Wang et al., 2024a) 7B AR 15M 512 0.75 0.60 7.76
ILLUME+ (Huang et al., 2025a) 7B AR 15M 512 0.75 0.60 7.76
TokenFlow (Qu et al., 2024) 7B AR 60M 256 0.65 0.65 7.78
Liquid (Wu et al., 2024b) 7B AR 30M 512 0.83 0.65 5.47
UniTok  (Ma et al., 2025) 7B AR 30M 256 0.85 0.67 7.46
SemHiTok(ours) 7B AR 15M 256 0.83 0.64 5.40
SemHiTok(ours) 7B AR 15M 384 0.83 0.66 5.70
Image Reconstruction. We present the reconstruction performance on the ImageNet-50k validation set in Tab.1. Notably, SemHiTok excels in reconstruction quality compared to the unified tokenizer, recording an impressive 1.16 rFID with $16\times$ downsampling ratio. While SemHiTok’s vocabulary size(approx 196k) appears larger than baselines, it is crucial to consider the effective representational capacity defined by the code shape. VILA-U(using RQ) and FQGAN(using Product Quantization) operate within a combinatorial search space($N^{D}$ or $N_{1}\times N_{2}$), resulting in a significantly larger effective capacity than SemHiTok’s linearly constrained structure($K\times m$). Furthermore, these approaches typically employ a denser code shape, whereas SemHiTok relies on a single unified hierarchical index. This indicates that our superior performance derives from the structured efficiency of SGHC rather than brute-force capacity expansion. Increasing the training 384 resolution led to a significant improvement in the rFID score, reaching 0.66. The results validate the effectiveness of SGHC design in modeling pixel feature space of the semantic code.
LLaVa-v1.5 Multimodal Understanding. To ensure a fair comparison, we conduct experiments to evaluate the multimodal understanding performance of existing open-source tokenizers and SemHiTok under the standard LLaVA-v1.5 setting. The results are as shown in Tab.2. LlamaGen’s performance is notably inferior, a consequence of its deficient pre-alignment with text. In contrast, unified image tokenizers show considerable advancements in understanding. However, due to their inherent hybrid structures or joint training strategies, a substantial disparity in understanding performance remains between prior discrete tokenizers and continuous representations. Notably, even though TokLIP uses a stronger base model (Qwen2.5-7B-Ins), its performance remains inferior to ours. SemHiTok achieves state-of-the-art results for discrete tokenizers, nearing the performance of continuous inputs such as SigLIP.
3.3 Unified MLLM
Multimodal Understanding. We evaluate the understanding performance of SemHiTok on diverse benchmarks in Table 4. Und means support multimodal understanding, Gen means support text-to-image generation. Among Und&Gen. Discrete, SemHiTok achieves state-of-the-art performance on most metrics, such as SEED, MMB, MME, and MME-P. Compared to other unified tokenizers (e.g., VILA-U, TokenFlow, UniToken), our models demonstrate significant advantages. Notably, our model surpasses expert-level models on key benchmarks, achieving 3.8 and 6.4 points higher than ShareGPT4V on MMMU and MMB, respectively. This bridges the gap between discrete visual tokens and continuous visual tokens in multimodal understanding tasks, strongly demonstrating the superiority and potential of our approach. Visualizations on understanding tasks are available in Supplementary Material 7.5.
Text-to-Image Generation To evaluate the text-to-image generation, we use GenAI-bench (Li et al., 2024a) and MJHQ30K (Li et al., 2024c) benchmark, and the results are shown in Tab4. For GenAI-bench, we use clip-flant5-xxl as the VQA score model to reflect the consistency between text descriptions and generated images. On this challenging benchmark, our model achieves competitive performance, closely matching Liquid, even though Liquid employs a generation-focused tokenizer. Furthermore, our model even outperformed some diffusion-based expert models, such as SDXL and SD v2. The strong results underscore the superior capability of our unified MLLM in complex text-to-image generation tasks. For MJHQ30K, we use the generation FID metric on generated images and high-quality images. On this benchmark, SemHiTok-256 attains 5.40 gFID, setting a new state-of-the-art in autoregressive image generation. In addition, we conduct further quantitative comparisons on GenEval and DPG, which are available in Supplementary Material 7.6. More quantitative analysis results and visualizations on generation tasks are available in Supplementary Material 7.7.
4 Ablation
4.1 Impact of Key Design.
Table 5: Impact of key design choices on reconstruction and multimodal understanding.
@@ 4 Ablation
In Tab.5, we validate the impact of our key design choices in SemHiTok: semantic codebook, semantic-guided hierarchical codebook(SGHC), Dual MLP, and Enhance Deocoder. For efficiency, we only test the understanding performance under LLaVA-v1.5, and train 40 epochs on ImageNet-1K for reconstruction tasks. We begin with the semantic codebook, which suffices for multimodal understanding but suffers from poor reconstruction. Incorporating SGHC enables pixel reconstruction and reduces rFID by 1.75, without noticeably affecting understanding performance. Introducing Dual-MLP further enhances multimodal understanding, even surpassing the semantic codebook alone, highlighting the effectiveness of multi-level feature modeling. Finally, a stronger pixel decoder brings additional improvements in reconstruction quality. figure 5 demonstrates the effects of various modules on reconstruction results. Compared to the semantic codebook, SGHC delivers more detailed reconstructions. Additionally, the enhanced decoder boosts performance on difficult samples. More reconstruction comparisons are available in Supplementary Material 7.8.
4.2 Impact of training strategy and structure.
We conduct comparative experiments under the same hyperparameters to investigate the impact of the training strategy and structure, as shown in Table 6. It can be observed that Exp 1’s performance is the lowest across both multimodal understanding and reconstruction tasks. After decoupling the architecture, Exp 2 achieves superior performance compared to Exp 1, yet its codebook usage remains at a low level. Furthermore, Exp 4 achieves marked gains in both efficacy and utilization over Exp 2, which underscores the importance of a phased training strategy for the proposed SemHiTok hierarchical architecture. Compared with Exp 3, Exp 4 achieves superior performance on both multimodal understanding and reconstruction tasks. This suggests that naively incorporating pixel features may negatively affect the alignment between image features and the LLM. In contrast, Exp 4 leverages pixel features as a complementary refinement to semantic features, thereby effectively bridging the gap between semantic and pixel representations.
4.3 More comparisons and ablation studies
In Supplementary Material 7.9, we attempt to provide quantitative evidence to further demonstrate pixel similarity within the semantic code. Meanwhile, to further illustrate how different hyperparameters in our method affect model performance, we also conduct ablation studies on Concat Type and Sub-Codebook Size in Supplementary Material 7.10, Semantic VQ Type, Codebook dim and Codebook size in Supplementary Material 7.11, and different $K\times m$ and more comparison with other large codebook reconstruction methods in Supplementary Material 7.12.
Table 6: Impact of training strategy and structure. DStruct: Using different learnable codebooks to model semantic and pixel information; DTrain: Adopt a phased optimization training strategy. Exp 1: used the same architecture as SDE and adopted a joint training strategy; Exp 2: using the same architecture as SemHiTok but with joint training; Exp 3: using two separately pretrained tokenizers (Llamagen and SigLIP) to independently extract semantic and pixel information.
Exp DStruct DTrain Multimodal understanding Reconstruction
@@ 4.3 More comparisons and ablation studies
1 58.0 1240.0 56.7 77.3 3.78 92.9%
2 57.8 1357.4 55.3 80.4 3.22 45.9%
3 58.7 1210.9 56.1 80.1 2.19 97.0%
4 SemHiTok 60.3 1355.8 62.9 82.6 1.42 93.7%
5 Related Work
5.1 Specialized Image Tokenizer
Tokenization for Autoregressive Generation. VQVAE(Van Den Oord et al., 2017) learns a discrete representation using a learnable codebook in auto-encoder architectures. VQGAN further improves the perceptual quality by using adversarial training(Goodfellow et al., 2014). improved the architecture with perceptual and adversarial-based losses for better reconstruction. This approach yields more precise and detailed image representations, significantly improving upon previous methodologies in image generation and processing. In recent literature, researchers are turning to efficient codebook structures(Shi et al., 2024; Zhang et al., 2023; Yu et al., 2021b; Bai et al., 2024) and better quantization methods(Zha et al., 2024; Yu et al., 2025) to improve generation performance and compression rates. IBQ(Shi et al., 2024) proposes the Index Backpropagation Quantization codebook update method, achieving stable training of large-scale codebooks. FQGAN(Bai et al., 2024) uses multiple codebooks and product quantization, where each codebook encodes a different type of feature. However, this work focuses only on image reconstruction and generation tasks, performing poorly on multimodal understanding tasks(Wu et al., 2024b; Liu et al., 2024a).
Tokenization for Understanding. In multimodal large language models (MLLMs)(Li et al., 2023b; 2022; Radford et al., 2021b; Liu et al., 2023a; Bai et al., 2023a; Chen et al., 2024c), researchers leverage CLIP(Radford et al., 2021b) and BLIP(Li et al., 2022) to extract visual characteristics that align with the language during its pre-training phase. Building upon many works (Liu et al., 2023a; Bai et al., 2023a; Chen et al., 2024c) have been collected and trained on high-quality datasets to achieve remarkable performance. LLaVA(Liu et al., 2023a) utilizes a vision encoder to align the vision inputs before LLMs. QwenVL(Bai et al., 2023a) and InterVL(Chen et al., 2024c) achieve better results through increased resolution, higher-quality datasets, etc. However, these text-aligned image encoders tend to focus on semantic information and ignore texture information, which is important for the generation task.
5.2 Unified Image Tokenizer
Numerous efforts have emerged to develop unified visual generation and understanding within one MLLM(Wang et al., 2024a; c; Xie et al., 2024; Dong et al., 2023; Ge et al., 2024; Sun et al., 2024b; Team, 2024; Wu et al., 2024d). There are two main lines to bridge the gap. Many workers (Dong et al., 2023; Ge et al., 2024; Sun et al., 2024b) combine diffusion models with LLMs for image generation. DREAMLLM(Dong et al., 2023) presents a unified framework that not only provides multimodal understanding but also creates multimodal content via diffusion models. Emu2(Sun et al., 2024b) trains a unified generative model using a diffusion-based decoder. These approaches inevitably increase model complexity and are not simple enough. Other workers (Team, 2024; Wu et al., 2024d; Xie et al., 2024; Wang et al., 2024c) adopt VQVAE-based encoders to convert images into discrete tokens. Chameleon(Team, 2024) and EMU3(Wang et al., 2024c) directly use VQGAN(Yu et al., 2021a), which is optimized by pixel reconstruction as the image tokenizer, while this method increases resource consumption during the pre-training stage and degrades multimodal understanding performance. VILA-U(Wu et al., 2024d) introduces a unified image tokenizer that incorporates a text-aligned branch within the VQGAN training paradigm. However, due to the gap between the semantic feature and the pixel feature, the joint optimization approach may lead to suboptimal solutions In contrast, our proposed SemHiTok can add the ability of extracting texture features without changing the discrete semantic features, and avoids the challenges brought by joint optimization.